# Yufei Bo, Shuo Shao, Meixia Tao

### v1

@@ I Introduction
II Overall Framework
Figure. 1 illustrates the overall framework of the DeepSCM scheme over a two-user Gaussian degraded broadcast channel. Without loss of generality, we assume that Receiver 2 has a higher channel SNR than Receiver 1. There is an observable source $\mathbf{X}$ associated with implicit hierarchical semantic information, namely the coarse-grained semantic information $\mathbf{S}_{1}$ and the fine-grained semantic information $\mathbf{S}_{2}$. Naturally, they form a Markov chain as $\mathbf{S}_{1}\to\mathbf{S}_{2}\to\mathbf{X}$. Following the conventional setup of semantic communications[37, 12], the receiver needs to recover both the semantic information and the observable source, which reflects the practical demands in many real-world scenarios where both humans and machines are involved in the task decision. Receiver 1 requires the observable source $\mathbf{X}$ and the coarse-grained semantic information $\mathbf{S}_{1}$, the recovery of which we respectively denote as $\hat{\mathbf{X}}_{1}$ and $\hat{\mathbf{S}}_{1}$. Meanwhile, as Receiver 2 has a larger channel capacity, it requires the observable source $\mathbf{X}$ and the fine-grained semantic information $\mathbf{S}_{2}$, the recovery of which we respectively denote as $\hat{\mathbf{X}}_{2}$ and $\hat{\mathbf{S}}_{2}$.
Specifically, in this paper we focus on image semantic communications for both image recovery and classification. Namely, the observable source $\mathbf{X}\in\mathbb{R}^{k}$ is image data, where $k$ represents the dimension of the images. The coarse-grained semantic information $\mathbf{S}_{1}\in\{1,2,...,L_{1}\}$ is the label of coarse image classification with $L_{1}$ classes. Similarly, the fine-grained semantic information $\mathbf{S}_{2}\in\{1,2,...,L_{2}\}$ is the label of fine image classification with $L_{2}$ classes. This hierarchical structure of the observable source and the semantics is derived from real-world scenarios. For instance, in wildlife monitoring, semantic information $\mathbf{S}_{1}$ can represent the general categories of animals, such as birds, bears, and so on, while semantic information $\mathbf{S}_{2}$ can represent the specific species within those categories, such as sparrows, cardinals, robins, black bears, brown bears, grizzly bears, etc.
The transmitter extracts semantic features in a hierarchical way, corresponding to the hierarchical sources intended for the two receivers. At the transmitter, there are two joint coding-modulation (JCM) blocks responsible for generating two constellation sequences in the superposition structure. Specifically, the first JCM block generates the inner constellation sequence $\mathbf{Y}_{1}\in\mathbb{C}^{n}$, which is intended for both receivers and carries the first level of semantic features to recover the observable source $\mathbf{X}$ and the coarse-grained semantic information $\mathbf{S}_{1}$. We use $n$ to denote the number of channel uses. Furthermore, the second JCM block generates the outer constellation sequence $\mathbf{Y}_{2}\in\mathbb{C}^{n}$, which is intended for Receiver 2 whose channel capacity is larger. The outer constellation sequence carries the additional semantic features, which together with $\mathbf{Y}_{1}$, recovers the observable source $\mathbf{X}$ and the fine-grained semantic information $\mathbf{S}_{2}$.
The first JCM block consists of a basic semantic encoder and a modulator. The basic semantic encoder $f_{\bm{\theta}_{1}}(\cdot)$ with parameters $\bm{\theta}_{1}$ extracts and encodes the basic semantic features from the observable source $\mathbf{X}$, outputting an $2n$-dimensional basic encoded feature vector $\mathbf{U}_{1}\in\mathbb{R}^{2n}$
$\mathbf{U}_{1}=f_{\bm{\theta}_{1}}(\mathbf{X}).$ (1)
Then, a probabilistic modulator with parameters $\bm{\alpha}_{1}$ generates $\mathbf{Y}_{1}$ from $\mathbf{U}_{1}$ by first learning the transition probability $p(\mathbf{y}_{1}|\mathbf{u}_{1},\bm{\alpha}_{1})$ then randomly sampling a sequence according to this transition probability as in our previous work [23].
The second JCM block consists of an enhancement semantic encoder, an LMMSE decorrelator and a modulator. The enhancement semantic encoder $f_{\bm{\theta}_{2}}(\cdot)$ with parameters $\bm{\theta}_{2}$ generates an enhanced encoded feature vector $\mathbf{U}_{2}\in\mathbb{R}^{2n}$ from $\mathbf{X}$, which contains the semantic features of $\mathbf{X}$ and $\mathbf{S}_{2}$
$\mathbf{U}_{2}=f_{\bm{\theta}_{2}}(\mathbf{X}).$ (2)
Considering the redundancy in $\mathbf{U}_{2}$ due to the hierarchical semantics, the LMMSE decorrelator projects $\mathbf{U}_{2}$ as the successive refinement vector $\mathbf{R}$ of the basic encoded feature vector $\mathbf{U}_{1}$, so that $\mathbf{R}$ and $\mathbf{U}_{1}$ can be uncorrelated. A detailed discussion on the LMMSE decorrelator will be further provided in Section III. Then, in a same manner that $\mathbf{Y}_{1}$ is generated, a probabilistic modulator with parameters $\bm{\alpha}_{2}$ learns the transition probability $p(\mathbf{y}_{2}|\mathbf{R},\bm{\alpha}_{2})$ then randomly samples $\mathbf{Y}_{2}$.
Figure 3: The NN framework of the JCM block. The output dimension of each layer is noted at the bottom.
In this paper, data transmission is carried out with $M$-QAM digital modulation by superposing $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$. Each element in $\mathbf{Y}_{1}$ takes values from an $M_{1}$-QAM constellation $\mathcal{C}_{1}=\{c_{1}^{I}+j\cdot c_{1}^{Q}\}$, and each element in $\mathbf{Y}_{2}$ takes values from an $M_{2}$-QAM constellation $\mathcal{C}_{2}=\{c_{2}^{I}+j\cdot c_{2}^{Q}\}$, where $M_{1}$ and $M_{2}$ satisfy $M_{1}\times M_{2}=M$. The output of the transmitter, which we denote as $\mathbf{Y}\in\mathbb{C}^{n}$, is formed by
$\mathbf{Y}=\sqrt{a}\mathbf{Y}_{1}+\sqrt{1-a}\mathbf{Y}_{2},$ (3)
where $a$ denotes the power allocation factor (PAF), and we set $a\in(0.5,1)$ so that the inner constellation is truly “inner”. Note that before superposition, we respectively scale $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ so that they meet an average transmit power constraint $P$, i.e., $\frac{\|\mathbf{Y}_{1}\|^{2}}{n}=\frac{\|\mathbf{Y}_{2}\|^{2}}{n}=P$. Receiver 1 receives a sequence $\mathbf{Z}_{1}=\mathbf{Y}+\bm{\epsilon}_{1}$, where $\bm{\epsilon}_{1}\sim\mathcal{CN}(0,\sigma^{2}_{1})$. Similarly, Receiver 2 receives a sequence $\mathbf{Z}_{2}=\mathbf{Y}+\bm{\epsilon}_{2}$ with $\bm{\epsilon}_{2}\sim\mathcal{CN}(0,\sigma^{2}_{2})$ and $\sigma_{1}>\sigma_{2}$. The channel condition of Receiver $i$ ($i=1,2$) is characterized by the channel SNR, which is defined as $\frac{P}{\sigma_{i}^{2}}$. Fig. 2 displays examples of superposed inner and outer constellations (which we henceforward call super-constellations). Fig. 2(a), (b), and (c) depict the super-constellation resulting from the superposition of two 4QAM constellations, denoted as the 4QAM×$\times$4QAM super-constellations. Fig. 2(d), (e), and (f) illustrate the super-constellation obtained by superposing a 4QAM inner constellation and a 16QAM outer constellation, denoted as the 4QAM×$\times$16QAM super-constellation. As a side note, the $M$-QAM digital modulation can be easily extended to other modulation schemes, such as $M$-PSK.
At each receiver, two decoders are deployed. The decoders at Receiver 1 are denoted as
$\mathbf{\hat{S}}_{1}$ $=g_{\bm{\psi}_{1}}(\mathbf{Z}_{1}),$ (4)
$\mathbf{\hat{X}}_{1}$ $=g_{\bm{\eta}_{1}}(\mathbf{Z}_{1}),$ (5)
where $\bm{\psi}_{1}$ parameterizes the decoder for classification and $\bm{\eta}_{1}$ parameterizes the decoder for image recovery. Similarly, the decoders at Receiver 2 are denoted as
$\mathbf{\hat{S}}_{2}$ $=g_{\bm{\psi}_{2}}(\mathbf{Z}_{2}),$ (6)
$\mathbf{\hat{X}}_{2}$ $=g_{\bm{\eta}_{2}}(\mathbf{Z}_{2}),$ (7)
with $\bm{\psi}_{2}$ and $\bm{\eta}_{2}$ respectively parameterizing the decoder for classification and image recovery.
III Transmitter Design
The core of our DeepSCM scheme lies in transmitter design, which aims at extracting uncorrelated semantic features and generating the inner and outer constellation sequences in the superposition structure. Therefore, this section is dedicated to describing the components of the transmitter as well as the training algorithm. We first discuss the basic building block of the transmitter, namely the JCM block. Then, in III-B, we explain the technical detail of the aforementioned LMMSE decorrelator. Finally, in III-C, we present the training algorithm of the whole system.
III-A Joint Coding-Modulation
Extending our previous work[23, 24] to a broadcast setting, we use two JCM blocks to enable the digital modulation of two encoded feature vectors. The NN framework of the first JCM block is shown in Fig. 3. The basic semantic encoder includes multiple Resnet blocks[18] to map an image $\mathbf{X}$ of dimension $W\times H\times C$ into the basic encoded feature vector $\mathbf{U}_{1}$ of dimension $2n$, where $W$, $H$ and $C$ respectively denote the width, height and channel of the image, and we have $k=W\times H\times C$.
We then modulate $\mathbf{U}_{1}$ to obtain $\mathbf{Y}_{1}$. The modulation process is learned as a probabilistic model to avoid the inherent nondifferentiability problem. A multi-layer perceptron (MLP) with a $\text{PReLu}(\cdot)$ activation function and a normalization layer outputs the transition probability, $p(\mathbf{y}_{1}|\mathbf{u}_{1},\bm{\alpha}_{1})$. Since $\mathbf{Y}_{1}$ is discretely distributed with $M_{1}^{n}$ possible values, the transition probability is a discrete probability distribution with $M_{1}^{n}$ categories. To simplify the learning, we model each element of $\mathbf{Y}_{1}$ to be conditionally independent. Therefore, the total number of probability categories to be learned for $\mathbf{Y}_{1}$ decreases to $n\cdot M_{1}$. That is, for each element of $\mathbf{Y}_{1}$, the MLP respectively outputs $M_{1}$ un-normalized probabilities, which is then normalized to be a probability distribution. Based on this probability distribution, the constellation symbol generator, using a differentiable sampling technique called the Gumbel-Softmax[38], samples a constellation symbol for this element.
The enhancement semantic encoder and the modulator in the second JCM block follow the same framework. However, different from the first JCM block, before modulating $\mathbf{U}_{2}$, we first decorrelate it with $\mathbf{U}_{1}$.
@@ III-A Joint Coding-Modulation
TABLE II: The NN architecture of the proposed scheme, where $n$ is the number of channel uses.
@@
Resnet Block $\times$ 3 2$\times$2$\times$2n
Average Pool + Flatten 2n
@@
Semantic Decoder for Classification at Receiver 1 Spinalnet Block $\times$ 4 + MLP 20
Semantic Decoder for Classification at Receiver 2 Spinalnet Block $\times$ 4 + MLP 100
Semantic Decoder for Image Recovery at Receiver 1 $\setminus$ Receiver 2 Conv + BatchNorm + ReLU 4$\times$4$\times$512
@@
III-B LMMSE Decorrelator
Due to the hierarchical relationship of the semantics, the enhanced encoded feature vector $\mathbf{U}_{2}$ is highly correlated with the basic encoded feature vector $\mathbf{U}_{1}$. Directly superposing them will cause redundancy in transmission and decrease the transmission efficiency.
To achieve maximum transmission efficiency, it is important that the encoded feature vectors being superposed should be independent. However, since independence between two variables is difficult to achieve, our goal is to remove from $\mathbf{U}_{2}$ the information that is highly correlated with $\mathbf{U}_{1}$ so that we can obtain a low-rate encoded feature vector that carries supplementary information from $\mathbf{U}_{1}$. We call this low-rate encoded feature vector the successive refinement vector $\mathbf{R}$ of $\mathbf{U}_{1}$. In other words, we want to find both a function $F(\cdot)$ and a vector $\mathbf{R}$, so that $\mathbf{U}_{2}$ can be written as
$\mathbf{U}_{2}=F(\mathbf{U}_{1})+\mathbf{R},$ (8)
where $\mathbf{R}$ should be low-rate since it carries less information. For simplicity, in this paper we choose a linear model for $F(\cdot)$. Hence, $\mathbf{U}_{2}$ can be written as
$\mathbf{U}_{2}$ $=F(\mathbf{U}_{1})+\mathbf{R}$
$=\mathbf{WU}_{1}+\mathbf{b}+\mathbf{R},$ (9)
where matrix $\mathbf{W}\in\mathbb{R}^{2n\times 2n}$ and vector $\mathbf{b}\in\mathbb{R}^{2n}$ are parameters to be determined so that a low-rate $\mathbf{R}$ can be obtained. However, it is hard to directly optimize the entropy of $\mathbf{R}$, given the challenge of estimating the entropy rate of a random variable. The following proposition presents an alternative approach to finding a low-rate $\mathbf{R}$.
Proposition 1 (Upperbound of the entropy)
The entropy of $\mathbf{R}$ is upperbounded by
$h(\mathbf{R})\leq n\log\frac{\pi e}{n}\mathbf{E}[||\mathbf{R}||_{2}^{2}].$ (10)
The proof can be found in Appendix A. As can be seen, the upperbound of the entropy of $\mathbf{R}$ is decided by the term $\mathbf{E}[||\mathbf{R}||_{2}^{2}]$. Therefore, minimizing the entropy of $\mathbf{R}$ can be translated into minimizing its expected $l_{2}$ norm. Expanding the term $\mathbf{E}[||\mathbf{R}||_{2}^{2}]$, we can get
In our DeepSCM scheme, we utilize a one-layer MLP for $F(\cdot)$ with input $\mathbf{U}_{1}$ and output ${\mathbf{W}}\mathbf{U}_{1}+\mathbf{b}$, where $\mathbf{W}$ and $\mathbf{b}$ respectively represent the weights and bias of the one-layer MLP. The optimization problem is thus established as
Notice that, the global optimal solution of this minimization problem is the LMMSE estimation of $\mathbf{U}_{2}$ given $\mathbf{U}_{1}$, where
[…]
Moreover, following the orthogonality principle of linear estimators [39, p.386], the global optimal $\mathbf{R}^{\ast}$ which serves as the estimation error, is uncorrelated with $\mathbf{U}_{1}$, i.e.,
$\mathbf{E}[\mathbf{U}_{1}\mathbf{R}^{\ast T}]=\mathbf{0}.$ (15)
In other words, the correlation between the $\mathbf{U}_{1}$ and $\mathbf{R}$ can indeed vanish with a sufficient number of training samples.
III-C Training Strategy
As can be seen, there are a total of nine NN modules used in the proposed scheme for different purposes. The intricate design of this framework makes it difficult to ensure the convergence of the network training. Therefore, we adopt a three-stage training strategy for our DeepSCM scheme. In the first stage, the encoding and decoding NNs for the generation of the inner constellation sequence are trained, namely the first JCM block and the decoders at Receiver 1. Then in the second stage, we further train the encoding and decoding NNs for the generation of the outer constellation sequence, including the second JCM block and the decoders at Receiver 2 while freezing other parameters. In the last training stage, the whole system is fine-tuned jointly to further improve its overall performance.
As a commonly used method[23, 13], we employ the cross entropy (CE) as the distortion measure function to evaluate the quality of classification, and the MSE to measure the distortion between the raw image and its recovery. Therefore, the loss function for the first training stage can be formulated as
$\mathcal{L}_{1}(\bm{\theta}_{1},\bm{\alpha}_{1},\bm{\psi}_{1},\bm{\eta}_{1})=CE(\mathbf{S}_{1},\mathbf{\hat{S}}_{1})+\lambda_{1}\cdot MSE(\mathbf{X},\mathbf{\hat{X}}_{1}),$ (16)
where $\lambda_{1}$ represents the hyperparameter to balance the two tasks, as used in many previous works[37, 12].
Similarly, the loss function of the second stage includes CE for fine classification, MSE for image recovery, and the objective in (12) as an additional regularizer, which can be written as
$\ \ \ \ \ \ \ \ \ \ \ \ \ \mathcal{L}_{2}(\bm{\theta}_{2},\bm{\alpha}_{2},\mathbf{W},\mathbf{b},\bm{\psi}_{2},\bm{\eta}_{2})=$
$CE(\mathbf{S}_{2},\mathbf{\hat{S}}_{2})+\lambda_{2}\cdot MSE(\mathbf{X},\mathbf{\hat{X}}_{2})+\lambda_{3}\cdot\|\mathbf{R}\|^{2}_{2},$ (17)
with $\lambda_{2}$ and $\lambda_{3}$ being two hyperparameters. Finally, the loss function in the fine-tuning stage is the combination of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, which is expressed as
$\mathcal{L}_{3}(\bm{\theta}_{1},\bm{\alpha}_{1},\bm{\psi}_{1},\bm{\eta}_{1},\bm{\theta}_{2},\bm{\alpha}_{2},\mathbf{W},\mathbf{b},\bm{\psi}_{2},\bm{\eta}_{2})=\mathcal{L}_{1}+\beta\cdot\mathcal{L}_{2},$ (18)
where $\beta$ is used to balance the importance of the two receivers.
IV Experiment Results
In this section, we use simulations to validate the advantages of our proposed DeepSCM scheme. We first present the experiment settings in Section IV-A. Then, we compare the performance of the DeepSCM scheme with benchmarks across various transmission rates and levels of channel disparity respectively in Section IV-B and IV-C. Furthermore, to establish guidelines for determining the PAF, we conduct experiments in Section IV-D examining its impacts on the training process and the overall system performance.
TABLE III: Detailed learning rate and number of training epochs for each stage.
Stage Number of Training Epochs Initial LR Scheduler

### v2

@@ III-A Joint Coding-Modulation
The second JCM block follows the same architecture as the first one, except that the enhanced EFV $\mathbf{U}_{2}$ should be decorrelated with $\mathbf{U}_{1}$ before modulation.
III-B LMMSE Decorrelator
Due to the hierarchical relationship of the semantics, the enhanced EFV $\mathbf{U}_{2}$ is highly correlated with the basic EFV $\mathbf{U}_{1}$. Directly superposing them will cause redundancy in transmission and decrease the transmission efficiency.
To achieve maximum transmission efficiency, it is important that the EFVs being superposed are independent. Thus, our goal is to remove from $\mathbf{U}_{2}$ the information that is dependent on $\mathbf{U}_{1}$ so that we can obtain a reduced-rate EFV that carries supplementary information from $\mathbf{U}_{1}$. We call this reduced-rate EFV the successive refinement vector of $\mathbf{U}_{1}$, denoted as $\mathbf{R}$. In other words, we want to find both a function $F(\cdot)$ and a vector $\mathbf{R}$, so that $\mathbf{U}_{2}$ can be written as
$\mathbf{U}_{2}=F(\mathbf{U}_{1})+\mathbf{R},$ (11)
where $\mathbf{R}$ should have smaller entropy rate since it carries less information. For simplicity, in this paper we choose a linear model for $F(\cdot)$. Hence, $\mathbf{U}_{2}$ can be written as
$\mathbf{U}_{2}$ $=\mathbf{WU}_{1}+\mathbf{b}+\mathbf{R},$ (12)
where $\mathbf{W}\in\mathbb{R}^{2n\times 2n}$ and $\mathbf{b}\in\mathbb{R}^{2n}$ are parameters to be determined so that a reduced-rate $\mathbf{R}$ can be obtained. However, it is hard to directly minimize the entropy of $\mathbf{R}$, given the challenge of estimating the entropy rate of a random variable. The following proposition presents an alternative approach to finding a low-rate $\mathbf{R}$.
Proposition 1
The entropy of $\mathbf{R}$ is upperbounded by
$h(\mathbf{R})\leq n\log\frac{\pi e}{n}\mathbf{E}[||\mathbf{R}||_{2}^{2}],$ (13)
where the equality holds if each element of $\mathbf{R}$ follows an i.i.d. Gaussian distribution with zero mean.
The proof can be found in Appendix A. As can be seen, the upperbound of the entropy of $\mathbf{R}$ is decided by $\mathbf{E}[||\mathbf{R}||_{2}^{2}]$, i.e., the expected $l_{2}$-norm of $\mathbf{R}$. Therefore, minimizing the entropy of $\mathbf{R}$ can be translated into minimizing $\mathbf{E}[||\mathbf{R}||_{2}^{2}]$, which is much more tractable. By using the linear model in (12), the term $\mathbf{E}[||\mathbf{R}||_{2}^{2}]$ can be expanded as
The optimization problem is thus established as
Notice that, problem (15) is essentially an LMMSE estimation problem of $\mathbf{U}_{2}$ given $\mathbf{U}_{1}$, whose global optimal solution is given by
$\mathbf{W}^{\ast}$ $=\textit{Cov}[\mathbf{U}_{2},\mathbf{U}_{1}]\textit{Var}[\mathbf{U}_{1}]^{-1},$ (16)
Following the orthogonality principle of linear estimators [40, p.386], the global optimal $\mathbf{R}^{\ast}$, which serves as the estimation error, is linearly uncorrelated with $\mathbf{U}_{1}$, i.e.,
$\mathbf{E}[\mathbf{U}_{1}\mathbf{R}^{\ast T}]=\mathbf{0}.$ (18)
The direct use of LMMSE estimation is computationally demanding since it requires the computation of the inversion of large matrices. As such, we utilize a one-layer MLP for $F(\cdot)$ with input $\mathbf{U}_{1}$ and output ${\mathbf{W}}\mathbf{U}_{1}+\mathbf{b}$, where $\mathbf{W}$ and $\mathbf{b}$ respectively represent the weights and bias of the MLP. The MLP-based LMMSE decorrelator is both computationally efficient and compatible with the mini-batch SGD algorithm.
1
2Stage 1, input: Image samples $\mathbf{X}$ at Tx, image samples $\mathbf{X}$ and coarse classification labels $\mathbf{S}_{1}$ at Rx1.
3while stop criterion is not met do
4       Forward pass
5       Tx computes and broadcasts $\mathbf{Y}$;
6       Rx1 receives $\mathbf{Z}_{1}$, and computes $g_{\bm{\eta}_{1}}(\mathbf{Z}_{1})\to\mathbf{\hat{X}}_{1}$, $g_{\bm{\psi}_{1}}(\mathbf{Z}_{1})\to\mathbf{\hat{S}}_{1}$, and the loss $\mathcal{L}_{1}$ by (19);
7       Backward pass
8       Rx1 trains $\bm{\psi}_{1}$, $\bm{\eta}_{1}$, then transmits the gradients of the first layer of $g_{\bm{\eta}_{1}}$ and $g_{\bm{\psi}_{1}}$ through an error-free link to Tx;
9       Tx trains $\bm{\theta}_{1}$, $\bm{\rho}_{1}$;
10
11Stage 2, input: Image samples $\mathbf{X}$ at Tx, image samples $\mathbf{X}$ and fine classification labels $\mathbf{S}_{2}$ at Rx2.
12while stop criterion is not met do
13       Forward pass
14       Tx computes and broadcasts $\mathbf{Y}$;
15       Rx2 receives $\mathbf{Z}_{2}$, and computes $g_{\bm{\eta}_{2}}(\mathbf{Z}_{2})\to\mathbf{\hat{X}}_{2}$, $g_{\bm{\psi}_{2}}(\mathbf{Z}_{2})\to\mathbf{\hat{S}}_{2}$, and the loss $\mathcal{L}_{2}$ by (20);
16       Backward pass
17       Rx2 trains $\bm{\psi}_{2}$, $\bm{\eta}_{2}$, and transmits the gradients of the first layer of $g_{\bm{\eta}_{2}}$ and $g_{\bm{\psi}_{2}}$ through an error-free link to Tx;
18       Tx trains $\bm{\theta}_{2}$, $\bm{\rho}_{2}$;
19
20Stage 3, input: Image samples $\mathbf{X}$ at Tx, image samples $\mathbf{X}$ and coarse classification labels $\mathbf{S}_{1}$ at Rx1, image samples $\mathbf{X}$ and fine classification labels $\mathbf{S}_{2}$ at Rx2.
21while stop criterion is not met do
22       Forward pass
23       Tx computes and broadcasts $\mathbf{Y}$;
24       Rx1 receives $\mathbf{Z}_{1}$, and computes $g_{\bm{\eta}_{1}}(\mathbf{Z}_{1})\to\mathbf{\hat{X}}_{1}$, $g_{\bm{\psi}_{1}}(\mathbf{Z}_{1})\to\mathbf{\hat{S}}_{1}$, and the loss $\mathcal{L}_{1}$ by (19);
25       Rx2 receives $\mathbf{Z}_{2}$, and computes $g_{\bm{\eta}_{2}}(\mathbf{Z}_{2})\to\mathbf{\hat{X}}_{2}$, $g_{\bm{\psi}_{2}}(\mathbf{Z}_{2})\to\mathbf{\hat{S}}_{2}$, and the loss $\mathcal{L}_{2}$ by (20);
26       Backward pass
27       Rx1 trains $\bm{\psi}_{1}$, $\bm{\eta}_{1}$, and transmits the gradients of the first layer of $g_{\bm{\eta}_{1}}$ and $g_{\bm{\psi}_{1}}$ through an error-free link to Tx;
28       Rx2 trains $\bm{\psi}_{2}$, $\bm{\eta}_{2}$, and transmits the gradients of the first layer of $g_{\bm{\eta}_{2}}$ and $g_{\bm{\psi}_{2}}$ through an error-free link to Tx;
29       Tx trains $\bm{\theta}_{2}$, $\bm{\rho}_{2}$ using gradient descent with $\mathcal{L}_{3}$ in (21), i.e., $\mathcal{L}_{1}+\beta\mathcal{L}_{2}$;
Output: trained parameters $\bm{\theta}_{1}$, $\bm{\rho}_{1}$, $\bm{\psi}_{1}$, $\bm{\eta}_{1}$, $\bm{\theta}_{2}$, $\bm{\rho}_{2}$, $\mathbf{W}$, $\mathbf{b}$, $\bm{\psi}_{2}$, $\bm{\eta}_{2}$.
Algorithm 1 DeepSCM training algorithm.
III-C Training Strategy
As can be seen from Fig. 1, there are a total of nine NN modules used in the proposed DeepSCM scheme for different purposes. The intricate design of this framework makes it difficult to ensure the convergence of the network training. Therefore, we adopt a three-stage training strategy. In the first stage, the encoding and decoding NNs for the generation of the inner constellation sequence are trained, namely the first JCM block and the decoders at Receiver 1. Then in the second stage, we further train the encoding and decoding NNs for the generation of the outer constellation sequence, including the second JCM block and the decoders at Receiver 2 while freezing other parameters. In the last training stage, the whole system is fine-tuned jointly to further improve its overall performance.
Specifically, the loss function for the first training stage is defined as
$\mathcal{L}_{1}(\bm{\theta}_{1},\bm{\rho}_{1},\bm{\psi}_{1},\bm{\eta}_{1})=D_{s}(\mathbf{S}_{1},\mathbf{\hat{S}}_{1})+\lambda_{1}\cdot D_{x}(\mathbf{X},\mathbf{\hat{X}}_{1}),$ (19)
where $D_{s}(\cdot,\cdot)$ denotes the distortion measure for semantic information recovery, $D_{x}(\cdot,\cdot)$ denotes the distortion measure for image recovery, and $\lambda_{1}$ represents the hyperparameter to balance the two distortions, as used in many previous works[38, 12]. In our work, we employ the commonly used cross entropy (CE) [25, 13] as the distortion measure for classification, i.e., $D_{s}(\mathbf{S}_{1},\mathbf{\hat{S}}_{1})=\text{CE}(\mathbf{S}_{1},\mathbf{\hat{S}}_{1})$. For image recovery, two distortion measures are considered. One is mean square error (MSE), which is used when the image recovery performance is evaluated by peak signal-to-noise ratio (PSNR). When the multi-scale structural similarity index (MS-SSIM) is used as the performance metric, the distortion measure is set to be $1-$ MS-SSIM.
Similarly, the loss function of the second stage can be defined as
$\ \ \ \ \ \ \ \ \ \mathcal{L}_{2}(\bm{\theta}_{2},\bm{\rho}_{2},\mathbf{W},\mathbf{b},\bm{\psi}_{2},\bm{\eta}_{2})=$
$D_{s}(\mathbf{S}_{2},\mathbf{\hat{S}}_{2})+\lambda_{2}\cdot D_{x}(\mathbf{X},\mathbf{\hat{X}}_{2})+\lambda_{3}\cdot\|\mathbf{R}\|^{2}_{2},$ (20)
where the objective in (15), i.e., the $l_{2}$-norm of the successive refinement vector $\mathbf{R}$, is included as an additional regularizer, and $\lambda_{2}$ and $\lambda_{3}$ are two hyperparameters. Finally, the loss function in the fine-tuning stage is the combination of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, expressed as
$\mathcal{L}_{3}(\bm{\theta}_{1},\bm{\rho}_{1},\bm{\psi}_{1},\bm{\eta}_{1},\bm{\theta}_{2},\bm{\rho}_{2},\mathbf{W},\mathbf{b},\bm{\psi}_{2},\bm{\eta}_{2})=\mathcal{L}_{1}+\beta\cdot\mathcal{L}_{2},$ (21)
where $\beta$ is used to balance the importance of the two receivers.
The overall training procedure is outlined in Algorithm 1. More specifically, during each training iteration, the transmitter randomly selects a source sample from the dataset, processes it through the NN, and broadcasts the output sequence to the receivers. The environmental condition is sampled to enable the training process. Each receiver then reconstructs the source sample and the semantic information, computes the loss, and backpropagates the gradients to update NN parameters using the SGD algorithm. During backpropagation, each receiver sends the gradients of its first layer through an error-free link to the transmitter, so that the transmitter can compute the gradients of its own NN parameters.
IV Experiment Results
In this section, we validate the advantages of DeepSCM via experimental simulation. We first present the experiment settings in Section IV-A. Then, we compare the performance of DeepSCM with benchmarks across various transmission rates and channel disparity levels respectively in Section IV-B and IV-C. The robustness of DeepSCM against variations in channel conditions is further demonstrated in Section IV-D. Additionally, we conduct experiments in Section IV-E examining the impact of PAF in the super-constellation. Finally, we provide discussion on the computation complexity of the proposed method in Section IV-F.
@@ IV Experiment Results
TABLE II: NN architecture of the proposed scheme, where $n$ is the number of channel uses.
@@
[…]
@@
Semantic Decoder for Classification at Rx1 Spinalnet Block $\times$ 4 + MLP 20
Semantic Decoder for Classification at Rx2 Spinalnet Block $\times$ 4 + MLP 100
Semantic Decoder for Image Recovery at Rx1 $\setminus$ Rx2 Conv + BatchNorm + ReLU 4$\times$4$\times$512
@@
TABLE III: Learning rate and training epochs for each stage.
Stage Training Epochs Initial LR Scheduler